A restaurant manager wants to order more glasses. The ratio of the number of water glasses to juice glasses is 15 to 2. How many juice glasses did the restaurant manager order if she ordered 180 water glasses?
Answer by TimothyLamb(4379)

w/j = 15/2
180/j = 15/2
j = 180/(15/2)
j = 180*(2/15)
j = 24
---
answer:
she ordered 24 juice glasses
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php